# RAPTOR: A Foundation Policy for Quadrotor Control

### v1

@@ 1 Introduction
Figure 1: (A) Motivation. Comparison of the adaptation capabilities of humans, contemporary RL-based control policies and our RAPTOR method. (B) The RAPTOR Method. Overview over all stages of the RAPTOR architecture.
@@
Bearing Vertical Take-Off and Landing (VTOL) and hovering capabilities, Multirotor Aerial Vehicles (MAVs) have become a valuable platform for many real-world applications like package delivery  (?), infrastructure inspection and maintenance  (?), or search and rescue  (?). In addition to the VTOL and hovering capabilities, MAVs can be built from cheap, mass-produced Commercial Off-The-Shelf (COTS) parts and be scaled from light (tens of grams) and centimeter-sized to heavy (multiple kilograms) and meter-sized. This broad range allows for easy customization of the mechanical and electrical design for each particular application.
@@
On the other hand, this variability also leads to challenges concerning the control of the platform. Changing the mechanical design and even simple modifications like switching the propeller or battery type often requires the retuning of the cascade of classical controllers that is still most widely used today. Analogously, modern, nonlinear control approaches, like Model Predictive Control (MPC) or Reinforcement Learning, heavily rely on accurate system models.
@@
RL policies have recently gained widespread popularity for quadrotor control  (?, ?, ?, ?). While overcoming many challenges, the resulting policies either overfit a single dynamics model or rely on domain randomization without adaptation. Domain randomization of the dynamics parameters  (?) is a powerful tool, but it forces fully Markovian (non-adaptive) policies to be conservative. Under domain randomization, the policy is incentivized to take conservative actions, e.g., in a state/observation that is critical for some of the quadrotors (e.g., due to thrust or angular acceleration limitations), the policy would have to take the action that saves these quadrotors, even if the current platform is agile enough to take the optimal action to continue the task.
@@
In addition to overfitting a particular system model, in quadrotor racing, it has become standard to also overfit to a particular trajectory (e.g., a certain racetrack)  (?, ?). This trend limits broad adoption of these methods for real-world use cases because even just changing the environment requires a full retraining of the policy.
@@
Humans, unlike these methods, which need to be retrained from scratch for each particular platform and/or application, can adapt few-shot to new systems. An illustrative example is driving cars (cf. Figure 1A and Movie S1). Initially, humans require extensive training to be able to control a car smoothly and robustly. But when they are driving a new, unseen car, they can adapt quickly. The steering, brake, and accelerator response might be quite different from the one they have experienced before, but they usually only require a handful of tries to adapt.
@@
With this work, we aim to devise a control policy that can adapt to unseen system dynamics using only minimal data, similar to humans. We are inspired by the recent progress on foundation models in the vision  (?) and language domain  (?). The main premises of foundation models are:
@@
Therefore, the goal is that our policy can quickly adapt to novel systems by interacting with it and using the context/sequence of high-frequency interactions to reason about the dynamics. This can also be seen as emergent implicit system identification. The objective is to control the quadrotor, i.e., based on the observations, output low-level motor commands that achieve an objective (e.g., controlling the position). Since the optimal outputs are dependent on the system dynamics, the (recurrent) policy has to learn to implicitly identify the unobserved/latent dynamics variables on the fly (figuratively and literally).
We refer to it as emergent implicit system identification because it only needs to infer the parts of the quadrotor dynamics parameters that are relevant to the input/output behavior of the system. These relevant parts can, e.g., be ratios like thrust-to-weight ratio, torque-to-inertia ratio, but also aggregates of motor delays, thrust curves, etc. We never train any neural network to explicitly reconstruct system parameters. The only training objective is performing well in terms of the reward function. This fulfills the premise 2) of foundation models.
@@
The main research questions are:
@@
    Does the policy generalize to unseen quadrotors that are 1) in-distribution and 2) out-of-distribution?
@@
We tackle the generalization question 4) by testing the policy on $9$ unseen but in-distribution (in terms of thrust-to-weight ratio, torque-to-inertia ratio, motor delays, thrust curves, etc.) quadrotors in the real world. We also study the out-of-distribution performance by testing the foundation policy (a) on a quadrotor with a flexible frame, (b) by installing four different propellers (2- and 3-blade), (c) by hitting it with a tool during flight, and (d) by testing it with a quadrotor that has thrust-to-weight ratio $2\times$ higher than the highest one experienced during training.
@@
We tackle the question 6) about the agility-adaptability trade-off by testing the resulting foundation policy on the task of tracking trajectories from quasi-static to highly dynamic.
Answering these research questions, we provide the following main contributions:
@@
    • A novel distillation method called Meta-Imitation Learning that condenses the behavior of $1000$ Markovian teacher policies into a single adaptive/non-Markovian student policy.
@@
While there have been many works on neural-network-based quadrotor control, most rely on lower-level controllers by, e.g., outputting collective thrust and body-rate setpoints and, hence, are not fully end-to-end  (?, ?, ?, ?, ?, ?). But recently there have also been works investigating full neural-network-based end-to-end control  (?, ?, ?, ?, ?). While these approaches have attained comparable performance to classical control schemes, the control policies are each highly optimized for a single quadrotor. Changing the quadrotor requires system identification to adjust the dynamics parameters of the simulator and a full retraining of the policy. This shortcoming is being tackled by the community right now, with works that investigate better adaptability and/or generalization of neural-network policies to multiple quadrotors. In particular,  (?) and  (?) are the most related works to our approach.
@@
In  (?), the authors train a single neural-network policy that can be deployed onto two different quadrotors with thrust-to-weight ratios of $\approx 5.8$ and $\approx 11.0$. The authors demonstrate impressive agility for racing through gates. The main difference to our approach is that their policy is Markovian/stateless, while ours is adaptive.
@@
Compared to  (?),  (?) is more related to our approach because the method intends to train a policy that can adapt to different quadrotors. In  (?), the authors show deployment of their adaptive policy to two relatively similar quadrotors with thrust-to-weight ratios of $3.23$ and $3.62$, respectively. The biggest differences to our work are that their approach is not end-to-end. A high-level controller that outputs collective thrusts and body rates is required. Their adaptive policy receives these CTBR setpoints as an input and is not concerned with rotational mechanics, etc. Our foundation policy is a full position controller and hence, in contrast to  (?), also understands the tilt required to build up linear velocity and execute translations as well as the angular velocities required to execute a particular tilt. Our policy covers these major non-linear transfer functions while their policy only covers the angular rate (and thrust control) that commonly is just implemented by a simple PD controller. Additionally, our policy architecture is simpler because it does not require training two encoders that map into the same latent space. Furthermore, our policy is vastly lighter, requiring only $2084$ parameters while theirs requires $114872$ parameters ($55\times$ larger), even though our policy covers more complex behavior. Because their policy is so computationally intensive, we cannot run it on any of the $10$ quadrotors we are using for testing without modifying the hardware to include a more powerful inference computer. Finally, the authors of  (?) do not publish their full training code, making it hard to exactly replicate their results. In contrast, we publish the full training and inference code in an easy and future-proof way. Additionally, readers can interact with the foundation policy through the web app at https://raptor.rl.tools.
2 Results
@@
2.1 Training
Figure 2: Training Results. (A) shows the pre-training learning curve, (B) shows the meta-imitation learning curve, (C) shows the Pareto frontier between performance and number of teachers, and (D) shows the Pareto frontier between performance and student/foundation policy size.
@@ 2.1 Training
In the following we describe the training results. The methods used to attain these results are described in the Materials and Methods Section 4. As described in Section 4 and Figure 1B, our training method separates into a pre-training and post-training phase.
2.2 Pre-Training
@@
In the pre-training phase, we train $1000$ teacher policies that each are specialized for a single quadrotor. The quadrotors differ in their dynamics parameters, which are sampled from the distribution described in Section 4.1.
@@
We find that the pre-training method described in Section 4.2.1 is remarkably robust and that all $1000$ RL training runs reliably converge to good policies, all of which can be used for the downstream Meta-Imitation Learning. This is surprising because RL training loops are known to be unstable and hence it is common to run the training with numerous seeds and then cherry-pick the best final policy  (?, ?, ?, ?, ?, ?). In contrast, we can use the same initial seed in all $1000$ cases, and we do not need to cherry-pick a different seed for each quadrotor. Note that our training runs are deterministic, i.e., a training run with the same seed always gives identical results (on the same computer). We also release the supplementary dataset Data S1 of $1000$ dynamics parameters and the $1000$ trained policies for each of them for further study.
@@
The aggregated learning curve resulting from the pre-training phase can be seen in 2A. We use the episode length as the metric here because it is more robust to the variation in quadrotor dynamics than the return. More agile quadrotors, for example, achieve much higher returns than less agile quadrotors, even if both are controlled by their respective optimal policy (with respect to the same reward function).
@@
We find that the policies already achieve good behavior after only $100000$ steps of training, which is signified by the majority of policies reaching episode lengths of $400$ steps or more. We observe that the main flight capabilities are attained in the first $100000$ steps, and that subsequent steps refine the behavior when starting in challenging initial conditions (large tilt, large linear velocity, position close to the termination boundary, …) as shown in the episode lengths converging to the maximum limit. We also observe that, in the later stages of the pre-training, the steady-state performance (which is not well expressed in the episode length) is also still improving.
@@
Since, in contrast to inference, the pre-training only incurs a one-time computational cost, we decide to train each teacher policy for $1$ M steps. Each training run takes $31$ m on a single core (consumer laptop, AMD Ryzen 9 7945HX, $16$ cores). Hence, the full pre-training takes about $34$ h on a single consumer laptop (cf. Section 4.2.1 on why and how the pre-training can be parallelized and sped up).
2.2.1 Meta-Imitation Learning
@@
None of these test-set quadrotors appear in the $1000$ pre-training quadrotors. We observe that the acquisition of good flying behavior follows a similar trajectory as a function of the epoch. For most quadrotors, the foundation policy achieves good performance early, but for the ”Flightmare” and especially the ”Large” model, we see continued improvement by training for longer. Eventually, after about $1000$ epochs, we observe convergence and good performance on all $7$ platforms, which span a large range of quadrotor dynamics, from agile to non-agile, lightweight to heavy, etc.
In Figure 2C, we study the performance of the resulting foundation policy when using the RAPTOR framework with different numbers of pre-training teachers/quadrotors. Here, the aggregate performance is shown across all $7$ test quadrotors, and we observe that, even for a low number of teachers, we see that the foundation policy is able to stabilize/control its position from various and challenging initial states without crashing.
As mentioned before, the episode length is an imperfect metric for the full performance of the policy, but, nevertheless, it is more suitable than the return, which strongly varies in scale, even just between the $7$ test-set quadrotors. With a training time of $1.9$ h (same AMD Ryzen 9 7945HX laptop), Meta-Imitation Learning is vastly less computationally intensive than the pre-training ($1.9$ h vs. $1000\times 31$ m).
@@
This small size allows the deployment to even the tiniest microcontrollers. Even on the most resource-constrained quadrotors, the resulting foundation policy requires $10\%$ of the available computational power, leaving plenty of capacity for state-estimation, communication and other duties of the flight controller.
@@
2.3 Emergent System Identification
@@
To answer the stated research questions we also study the behavior of the foundation policy resulting from Meta-Imitation Learning. We find that it indeed exhibits in-context learning as can be seen in Figure 3. Here we start a quadrotor in a state where it is displaced by $2$ m from the target position in x and y as well as pitched backwards by $90^{\circ}$ (no linear or angular velocity). First of all, we can see that the policy successfully manages to recover from this initial state and to reduce the distances towards the target state in position, orientation, linear and angular velocity.
Additionally, we plot the trajectory of the hidden state and investigate if it learns something about the dynamics of the quadrotor it is interacting with. We know the ground-truth dynamics parameters of the $1000$ quadrotors used for pre-training, hence we can train a linear probe  (?) to predict, e.g., the thrust-to-weight ratio. Linear probing has become a standard test to evaluate if foundation models learn good representations  (?, ?, ?).
As can be seen from the regression plot in Figure 3, just a linear probe can predict the thrust-to-weight ratio very well. We use an $80\%$-$20\%$ train-test split and the linear model achieves a Mean-Squared Error (MSE) of $0.047$ and an $R^{2}$ of $0.949$, significantly reducing the a priori uncertainty about the thrust-to-weight ratio. This shows that Meta-Imitation Learning leads to emergent implicit system identification in the latent space of the foundation policy.
While quickly going into the right range, we can also see that the estimate is improving over time, showing the in-context learning of the policy.
2.4 Deployment
@@
To show the robustness and adaptability of the foundation policy that results from applying the RAPTOR framework, we deploy the foundation policy onto $10$ different real quadrotors and $2$ different simulators (cf. Figure 4) showing Simulation-to-Reality (Sim2Real) and Simulation-to-Simulation (Sim2Sim) transfer. We aim to strain the adaptation capabilities of the foundation policy as much as possible by testing across a wide range of parameters:
  1. 1.
    A quantitatively wide range of parameters:
    • Weight: $31.9$ g - $2.4$ kg
    • Size: $65$ mm - $500$ mm
    • Thrust-to-weight: $\approx 1.75$ - $12$
  2. 2.
    A qualitatively diverse set of features:
    • Flight controller: PX4, Betaflight, Crazyflie, M5StampFly
    • State estimator: EKF, Mahony, Madgwick
    • Motor type: brushed and brushless
    • Flexible frame
    • Mixing two- and three-blade propellers
Many of these quantities are (far) out-of-distribution, like a thrust-to-weight ratio of $12$ ($\leq 5$ in training), a flexible frame (only rigid during training), and observations from a state estimator (Ground Truth (GT) during training).
This shows that our proposed RAPTOR framework actually produces a policy that not only generalizes to quadrotors that are in the training distribution (cf. Section 4.1) but also out-of-distribution (OOD).
Contrary to popular belief, and supporting the results in  (?), we find that the simulation-to-reality transfer of end-to-end neural network policies is actually not hampered by the mismatch in the dynamics model, especially because the parameters can be relatively easily and accurately determined using  (?). We find that qualitative difference matters much more. Especially for stateful policies (like the RNN used in the RAPTOR foundation policy), implementation details in the firmware that lead to delays and other artifacts have a strong impact on the simulation-to-reality deployment.
We find that the foundation policy works robustly on all platforms, but we also observe low-frequency z-axis oscillations in some of the non-EKF-based quadrotors. In the case of the Mahony  (?) and Madgwick  (?) filters, only the orientation is estimated by the filter and the velocity is directly fed from the motion capturing system.
We hypothesize that this leads to the z-axis oscillations due communication delays. We can reproduce the z-axis oscillations in simulation across quadrotors of different dynamics parameters by inducing a linear velocity delay of 103010-30 ms.
@@ 2.4.1 Trajectory Tracking
Figure 5 and Movie S2 show trajectory tracking of a Lissajous-based figure-eight trajectory at different intervals. We test tracking the $10$ s and $5.5$ s trajectories using the foundation policy on all $12$ quadrotors. Each of the shown trajectories constitutes $5$ consecutive full loops with an initial linear ramp-up time of $1$ s (from hovering). We find that the trajectory tracking performance is good in all real-world cases.
As can be seen from the Root Mean Square Error (RMSE) values in the bottom table, the tracking performance is in line with state-of-the-art policies (without trajectory lookahead, like the foundation policy) that are trained for a single quadrotor  (?). The dedicated policy that has been specifically trained for deployment on a Crazyflie in  (?) reaches an RMSE tracking error of $0.17$ m and $0.15$ m (with and without the z-axis) for the identical $5.5$ s figure-eight trajectory. The foundation policy resulting from the RAPTOR framework reaches $0.19$ m and $0.19$ m, respectively, on the same platform. While the tracking error of the foundation policy is slightly elevated compared to the dedicated policy, through online adaptation, the foundation policy can attain a similar performance on a plethora of other quadrotors as well.
@@
The third cluster of trajectories in Figure 5 shows the fastest trajectory that we tested for each quadrotor. This does not necessarily mean that this is the fastest trajectory supported by the foundation policy, because, to avoid crashes, we did not push all quadrotors (especially the larger ones) beyond their limits. For the Hummingbird, Crazyflie, M5StampFly, Crazyflie Brushless, and Meteor75 Pro, we did push them beyond their (or the foundation policy’s) limits and find that the shown trajectories are the most agile ones that can still be tracked with decent accuracy. When pushing, e.g., the Hummingbird or the Meteor75 Pro further, they still remain stable, but they overshoot so much that the figure-eight becomes barely recognizable.

### v2

@@ Abstract
Figure 1: (A) Motivation. Comparison of the adaptation capabilities of humans, contemporary RL-based control policies, and our RAPTOR method. (B) The RAPTOR Method. Overview of all stages of the RAPTOR architecture.
Introduction
@@
Bearing Vertical Take-Off and Landing (VTOL) and hovering capabilities, Multirotor Aerial Vehicles (MAVs) have become a valuable platform for many real-world applications like package delivery (?), infrastructure inspection and maintenance (?), or search and rescue (?). In addition to the VTOL and hovering capabilities, MAVs can be built from cheap, mass-produced Commercial Off-The-Shelf (COTS) parts and can be scaled from light (tens of grams) and centimeter-sized to heavy (multiple kilograms) and meter-sized. This broad range allows for easy customization of the mechanical and electrical design for each particular application.
@@
On the other hand, this variability also leads to challenges concerning the control of the platform. Changing the mechanical design and even simple modifications like switching the propeller or battery type often require the retuning of the cascade of classical controllers that is still most widely used today. Analogously, modern, nonlinear control approaches, like Model Predictive Control (MPC) or Reinforcement Learning, heavily rely on accurate system models.
@@
Reinforcement Learning (RL) policies have recently gained widespread popularity for quadrotor control (?, ?, ?, ?). Although overcoming many challenges, the resulting policies either overfit a single dynamics model or rely on domain randomization without adaptation. Domain randomization of the dynamics parameters (?) is a powerful tool, but it forces fully Markovian (non-adaptive) policies to be conservative. Under domain randomization, the policy is incentivized to take conservative actions, for example, in a state/observation that is critical for some of the quadrotors (for example, due to thrust or angular acceleration limitations), the policy would have to take the action that saves these quadrotors, even if the current platform is agile enough to take the optimal action to continue the task.
@@
In addition to overfitting a particular system model, in quadrotor racing, it has become standard to also overfit to a particular trajectory (such as a certain racetrack) (?, ?). This trend limits broad adoption of these methods for real-world use cases because even just changing the environment requires a full retraining of the policy.
@@
Humans, unlike these methods, which need to be retrained from scratch for each particular platform and/or application, can adapt few-shot to new systems. An illustrative example is driving cars (see Figure 1A and Movie S1). Initially, humans require extensive training to be able to control a car smoothly and robustly. But when they are driving a new, unseen car, they can adapt quickly. The steering, brake, and accelerator response might be quite different from the ones they have experienced before, but they usually only require a handful of tries to adapt.
@@
With this work, we aim to devise a control policy that can adapt to unseen system dynamics using only minimal data, similar to humans. We are inspired by the recent progress on foundation models in the vision (?) and language domains (?). The two main premises of foundation models are:
@@
Therefore, the goal is that our policy can quickly adapt to novel systems by interacting with them and using the context/sequence of high-frequency interactions to reason about the dynamics. This can also be seen as emergent implicit system identification. The objective is to control the quadrotor, that is, based on the observations, output low-level motor commands that achieve an objective (such as controlling the position). Since the optimal outputs are dependent on the system dynamics, the (recurrent) policy has to learn to implicitly identify the unobserved/latent dynamics variables on the fly (figuratively and literally).
We refer to this as emergent implicit system identification because it only needs to infer the parts of the quadrotor dynamics parameters that are relevant to the input/output behavior of the system. These relevant parts can, for example, be ratios like thrust-to-weight ratio, torque-to-inertia ratio, but also aggregates of motor delays, thrust curves, and so on. We never train any neural network to explicitly reconstruct system parameters. The only training objective is performing well in terms of the reward function. This fulfills the premise 2) of foundation models.
@@
The six main research questions are:
@@
    Does the policy generalize to unseen quadrotors that are in-distribution and out-of-distribution?
@@
We tackle the generalization question 4) by testing the policy on $9$ unseen but in-distribution (in terms of thrust-to-weight ratio, torque-to-inertia ratio, motor delays, and thrust curves) quadrotors in the real world. We also study the out-of-distribution performance by testing the foundation policy (a) on a quadrotor with a flexible frame (making it a total of 10 real quadrotors), (b) by installing four different propellers (2- and 3-blade), (c) by hitting it with a tool during flight, and (d) by testing it with a quadrotor that has thrust-to-weight ratio $2\times$ higher than the highest one experienced during training.
@@
We tackle question 6) about the agility-adaptability trade-off by testing the resulting foundation policy on the task of tracking trajectories from quasi-static to highly dynamic.
Answering these research questions, we provide the following five main contributions:
  1. 1.
@@
    • Our proposed distillation method called Meta-Imitation Learning that condenses the behavior of $1000$ Markovian teacher policies into a single adaptive/non-Markovian student policy.
@@
Although there have been many works on neural-network-based quadrotor control, most rely on lower-level controllers by, for example, outputting collective thrust and body-rate (CTBR) setpoints and, hence, are not fully end-to-end (?, ?, ?, ?, ?, ?). But recently there have also been works investigating full neural-network-based end-to-end control (?, ?, ?, ?, ?). Although these approaches have attained comparable performance to classical control schemes, the control policies are each highly optimized for a single quadrotor. Changing the quadrotor requires system identification to adjust the dynamics parameters of the simulator and a full retraining of the policy. This shortcoming is being tackled by the community right now, with works that investigate better adaptability and/or generalization of neural-network policies to multiple quadrotors. In particular, (?) and (?) are the most related works to our approach.
@@
In (?), the authors train a single neural-network policy that can be deployed onto two different quadrotors with thrust-to-weight ratios of $\approx 5.8$ and $\approx 11.0$, respectively. The authors demonstrate impressive agility for racing through gates. The main difference to our approach is that their policy is Markovian/stateless, whereas ours is adaptive.
@@
Compared to (?), (?) is more related to our approach because the method intends to train a policy that can adapt to different quadrotors. In (?), the authors show deployment of their adaptive policy to two relatively similar quadrotors with thrust-to-weight ratios of $3.23$ and $3.62$, respectively. The biggest differences to our work are that their approach is not end-to-end. A high-level controller that outputs collective thrusts and body-rates (CTBR) is required. Their adaptive policy receives these CTBR setpoints as an input and is not concerned with rotational mechanics. Our foundation policy is a full position controller and hence, in contrast to (?), also understands the tilt required to build up linear velocity and execute translations as well as the angular velocities required to execute a particular tilt. Our policy covers these major non-linear transfer functions whereas their policy only covers the angular rate (and thrust control) that is commonly implemented by a simple PD controller. Additionally, our policy architecture is simpler because it does not require training two encoders that map into the same latent space. Furthermore, our policy is vastly lighter, requiring only $2084$ parameters whereas theirs requires $114872$ parameters ($55\times$ larger), even though our policy covers more complex behavior. Because their policy is so computationally intensive, we cannot run it on any of the $10$ quadrotors we are using for testing without modifying the hardware to include a more powerful inference computer. Finally, the authors of (?) do not publish their full training code, making it hard to exactly replicate their results. In contrast, we publish the full training and inference code in an easy and future-proof way. Additionally, readers can interact with the foundation policy through the web app at https://raptor.rl.tools.
Beyond quadrotor control, distilling the behavior of a privileged teacher into a partially observable student has become increasingly popular for robotic control tasks. In (?), this idea is referred to as “Learning by Cheating”, and a policy for driving cars is distilled from a teacher that is based on imitation learning with dense states from expert demonstrations into a student that only receives visual inputs. Compared to our approach, the inference is not performed by integrating information over time, but by extracting the remaining information after a lossy projection (ego-view). In (?) and (?), the authors distill privileged teachers that can observe details about the terrain into student policies that only receive proprioceptive observations. Like in (?), the latter uses an environment encoder and shares the control head between teacher and student. Like in our work, the students are sequence models. Although distilling a privileged teacher into a partially observable student appears like a simple concept at a high level, there are many nuances and design decisions, many of which have been studied recently in (?). The authors show the effectiveness of a GRU-based adaptive policy and thoroughly ablate many important aspects, but only demonstrate it using a cartpole, which is a relatively simple system compared to the quadrupeds and quadrotors used in this and the other referenced works. In contrast to all other works, we do not restrict our method to a single teacher. This circumvents well-known challenges in multi-task RL, where “negative transfer” is often observed (?, ?), and additionally makes the pre-training stage embarrassingly parallel, which suits modern hardware well.
Results
@@
Training
@@
In the following, we describe the training results. The methods used to attain these results are described in the Materials and Methods section. As described in the Materials and Methods section and Figure 1B, our training method separates into a pre-training phase and a post-training phase.
Pre-Training
@@
In the pre-training phase, we train $1000$ teacher policies that are each specialized for a single quadrotor. The quadrotors differ in their dynamics parameters, which are sampled from the distribution described in the Domain Randomization subsection.
@@
We find that the pre-training method described in the Pre-Training subsection of the Materials and Methods section is remarkably robust and that all $1000$ RL training runs reliably converge to good policies, all of which can be used for the downstream Meta-Imitation Learning. This is surprising because RL training loops are known to be unstable and hence it is common to run the training with numerous seeds and then cherry-pick the best final policy (?, ?, ?, ?, ?, ?).
In contrast, we can use the same initial seed in all $1000$ cases, and we do not need to cherry-pick a different seed for each quadrotor. Note that our training runs are deterministic, that is, a training run with the same seed always gives identical results (on the same computer). We attribute this stability to a combination of three main factors:
  1. 1.
    Overparameterization: Since we do not need to deploy the teacher policies onto constrained hardware, we can use much larger ($\approx 55\times$) neural networks. This overparameterization is known to improve the loss landscape and stabilize training (?).
  2. 2.
    Reward Function: We tuned the reward function for reliable training, particularly the survival bonus and the termination penalty.
  3. 3.
    Off-Policy RL: We use an off-policy RL algorithm (SAC) because our experience matches the literature in that on-policy methods like PPO are usually more unstable and prone to local minima (?). We find off-policy methods to be particularly robust when the replay buffer retains all interaction steps and when combined with the previous two points.
If these conditions are met, it is just a matter of training long enough to see all training runs converge to a good policy (see Figure 2A). We also release the supplementary dataset Data S1 of $1000$ dynamics parameters and the $1000$ trained policies for each of them for further study.
Figure 2: Training Results. (A) shows the pre-training learning curve, (B) shows the meta-imitation learning curve where the policy is evaluated using a validation set of $7$ quadrotors that are not seen during training, (C) shows the Pareto frontier between performance and number of teachers, and (D) shows the Pareto frontier between performance and student/foundation policy size.
@@
The aggregated learning curve resulting from the pre-training phase can be seen in Figure 2A. We use the episode length as the metric here because it is more robust to the variation in quadrotor dynamics than the return. More agile quadrotors, for example, achieve much higher returns than less agile quadrotors, even if both are controlled by their respective optimal policy (with respect to the same reward function).
@@
We find that the policies already achieve good behavior after only $100000$ steps of training, which is signified by the majority of policies reaching episode lengths of $400$ steps or more. We observe that the main flight capabilities are attained in the first $100000$ steps, and that subsequent steps refine the behavior when starting in challenging initial conditions (large tilt, large linear velocity, position close to the termination boundary, etc.) as shown in the episode lengths converging to the maximum limit. We also observe that, in the later stages of the pre-training, the steady-state performance (which is not well expressed in the episode length) is also still improving.
@@
Since, in contrast to inference, the pre-training only incurs a one-time computational cost, we decide to train each teacher policy for $1$ M steps. Each training run takes $31$ min on a single core (consumer laptop, AMD Ryzen 9 7945HX, $16$ cores). Hence, the full pre-training takes about $34$ h on a single consumer laptop. Please refer to the Pre-Training subsection in the Materials and Methods section for details on why and how the pre-training can be parallelized and sped up.
Meta-Imitation Learning
@@
None of these test-set quadrotors appear in the $1000$ pre-training quadrotors. We observe that the acquisition of good flying behavior follows a similar trajectory as a function of the epoch. For most quadrotors, the foundation policy achieves good performance early, but for the ”Flightmare” and especially the ”Large” model, we see continued improvement by training for longer. Eventually, after about $1000$ epochs, we observe convergence and good performance on all $7$ platforms, which span a large range of quadrotor dynamics, from agile to non-agile, lightweight to heavy, and more.
In Figure 2C, we study the performance of the resulting foundation policy when using the RAPTOR framework with different numbers of pre-training teachers/quadrotors. Here, the aggregate performance is shown across all $7$ test quadrotors, and we observe that, even for a low number of teachers, the foundation policy is able to stabilize/control its position from various and challenging initial states without crashing.
As mentioned before, the episode length is an imperfect metric for the full performance of the policy, but, nevertheless, it is more suitable than the return, which strongly varies in scale, even just between the $7$ test-set quadrotors. With a training time of $1.9$ h (same AMD Ryzen 9 7945HX laptop), Meta-Imitation Learning is vastly less computationally intensive than the pre-training ($1.9$ h vs. $1000\times 31$ min).
@@
This small size allows the deployment to even the tiniest microcontrollers. Even on the most resource-constrained quadrotors, the resulting foundation policy requires $10\%$ of the available computational power, leaving plenty of capacity for state-estimation, communication, and other duties of the flight controller.
@@
Emergent System Identification
To answer the stated research questions, we also study the behavior of the foundation policy resulting from Meta-Imitation Learning. We find that it indeed exhibits in-context learning as can be seen in Figure 3. Here we start a quadrotor in a state where it is displaced by $2$ m from the target position in $x$ and $y$ as well as pitched backwards by $90^{\circ}$ (no linear or angular velocity). First of all, we can see that the policy successfully manages to recover from this initial state and to reduce the distances towards the target state in position, orientation, linear and angular velocity.
@@
Additionally, we plot the trajectory of the hidden state and investigate if it learns something about the dynamics of the quadrotor it is interacting with. We know the ground-truth dynamics parameters of the $1000$ quadrotors used for pre-training, hence we can train a linear probe (?) to predict, for example, the thrust-to-weight ratio. Linear probing has become a standard test to evaluate if foundation models learn good representations (?, ?, ?).
As can be seen from the regression plot in Figure 3, just a linear probe can predict the thrust-to-weight ratio very well. For this experiment, we collect $100$ new episodes (neither seen during Pre-Training nor Meta-Imitation Learning) from random initial states for each of the $1000$ sampled quadrotors. We use an $80\%$-$20\%$ train-test split and the linear model achieves a Mean-Squared Error (MSE) of $0.047$ and an $R^{2}$ of $0.949$, substantially reducing the a priori uncertainty about the thrust-to-weight ratio. This shows that Meta-Imitation Learning leads to emergent implicit system identification in the latent space of the foundation policy.
Although quickly going into the right range, we can also see that the estimate is improving over time, showing the in-context learning of the policy.
Deployment
To show the robustness and adaptability of the foundation policy that results from applying the RAPTOR framework, we deploy the foundation policy onto $10$ different real quadrotors and $2$ different simulators (see Figure 4) showing Simulation-to-Reality (Sim2Real) and Simulation-to-Simulation (Sim2Sim) transfer. We aim to strain the adaptation capabilities of the foundation policy as much as possible by testing across a wide range of parameters:
  1. 1.
    A quantitatively wide range of parameters: weight ($31.9$ g - $2.4$ kg), size ($65$ mm - $500$ mm), and thrust-to-weight ($\approx 1.75$ - $12$).
  2. 2.
    A qualitatively diverse set of features: flight controller (PX4, Betaflight, Crazyflie, M5StampFly), state estimator (EKF, Mahony, Madgwick), motor type (brushed and brushless), flexible frame, and mixing two- and three-blade propellers.
@@
Many of these quantities are (far) out-of-distribution, like a thrust-to-weight ratio of $12$ ($\leq 5$ in training), a flexible frame (only rigid during training), and observations from a state estimator (Ground Truth during training).
This shows that our proposed RAPTOR framework actually produces a policy that not only generalizes to quadrotors that are in the training distribution (see Domain Randomization subsection) but also out-of-distribution (OOD).
Contrary to popular belief, and supporting the results in (?), we find that the simulation-to-reality transfer of end-to-end neural network policies is actually not hampered by the mismatch in the dynamics model, especially because the parameters can be relatively easily and accurately determined using (?). We find that qualitative differences matter much more. Especially for stateful policies (like the RNN used in the RAPTOR foundation policy), implementation details in the firmware that lead to delays and other artifacts have a strong influence on the simulation-to-reality deployment.
We find that the foundation policy works robustly on all platforms, but we also observe low-frequency z-axis oscillations in some of the non-EKF-based quadrotors. In the case of the Mahony (?) and Madgwick (?) filters, only the orientation is estimated by the filter and the velocity is directly fed from the motion capture system.
We hypothesize that this leads to the z-axis oscillations due to communication delays. We can reproduce the z-axis oscillations in simulation across quadrotors of different dynamics parameters by inducing a linear velocity delay of 103010-30 ms.
@@
Figure 5 and Movie S2 show trajectory tracking of a Lissajous-based figure-eight trajectory at different intervals. We test tracking the $10$ s and $5.5$ s trajectories using the foundation policy on all $12$ quadrotors. Each of the shown trajectories constitutes $5$ consecutive full loops with an initial linear ramp-up time of $1$ s (from hovering).
Figure 5 shows consistent real-world trajectory-tracking performance across repetitions. The Root Mean Square Error (RMSE) ranges between $0.07$ m and $0.19$ m (mean: $0.11$ m, standard deviation: $0.04$ m) for the $10$ s figure-eight trajectory and between $0.17$ m and $0.29$ m (mean: $0.20$ m, standard deviation: $0.04$ m) for the $5.5$ s figure-eight trajectory.
The tracking performance is in line with state-of-the-art policies (without trajectory lookahead, like the foundation policy) that are trained for a single quadrotor (?). The dedicated policy that has been specifically trained for deployment on a Crazyflie in (?) reaches an RMSE tracking error of $0.17$ m and $0.15$ m (with and without the z-axis) for the identical $5.5$ s figure-eight trajectory. The foundation policy resulting from the RAPTOR framework reaches $0.19$ m and $0.19$ m, respectively, on the same platform. Although the tracking error of the foundation policy is slightly elevated compared to the dedicated policy, through online adaptation, the foundation policy can attain a similar performance on a plethora of other quadrotors as well. Note that our policy still performs better than the two other neural-network-based baselines evaluated in (?) which attain $0.23$ m / $0.21$ m (?) and $0.25$ m / $0.24$ m (?), respectively.
@@
The third cluster of trajectories in Figure 5 shows the fastest trajectory that we tested for each quadrotor. This does not necessarily mean that this is the fastest trajectory supported by the foundation policy, because, to avoid crashes, we did not push all quadrotors (especially the larger ones) beyond their limits. For the Hummingbird, Crazyflie, M5StampFly, Crazyflie Brushless, and Meteor75 Pro, we did push them beyond their (or the foundation policy’s) limits and find that the shown trajectories are the most agile ones that can still be tracked with decent accuracy. When pushing, for example, the Hummingbird or the Meteor75 Pro further, they still remain stable, but they overshoot so much that the figure-eight becomes barely recognizable.